# Reinforced MLLM: A Survey on RL-Based Reasoning in Multimodal Large Language Models

### v1

@@ Abstract.
The integration of reinforcement learning (RL) into the reasoning capabilities of Multimodal Large Language Models (MLLMs) has rapidly emerged as a transformative research direction. While MLLMs significantly extend Large Language Models (LLMs) to handle diverse modalities such as vision, audio, and video, enabling robust reasoning across multimodal inputs remains a major challenge. This survey systematically reviews recent advances in RL-based reasoning for MLLMs, covering key algorithmic designs, reward mechanism innovations, and practical applications. We highlight two main RL paradigms—value-free and value-based methods—and analyze how RL enhances reasoning abilities by optimizing reasoning trajectories and aligning multimodal information. Furthermore, we provide an extensive overview of benchmark datasets, evaluation protocols, and existing limitations, and propose future research directions to address current bottlenecks such as sparse rewards, inefficient cross-modal reasoning, and real-world deployment constraints. Our goal is to offer a comprehensive and structured guide to researchers interested in advancing RL-based reasoning in the multimodal era.
@@ 1. Introduction
The emergence of large language models (LLMs) (Yang et al., 2024; GLM et al., 2024; Grattafiori et al., 2024; Abdin et al., 2024; Team et al., 2023) has ushered in an unprecedented new era for the field of artificial intelligence, showcasing exceptional capabilities such as instruction following and few-shot learning (Brown et al., 2020). However, achieving human-level intelligence requires not only surpassing basic perceptual abilities but also developing sophisticated cognitive reasoning skills that enable iterative reasoning through contextual understanding and self-correction. Inspired by this, in-context learning (ICL) techniques (Wei et al., 2021, 2022a; Wu et al., 2022) have empowered LLMs with step-by-step reasoning abilities, commonly referred to as the Chain-of-Thought (CoT) reasoning mechanism (Wei et al., 2022b; Zhang et al., 2022; Wang and Zhou, 2024; Besta et al., 2024). OpenAI’s o1 model (Jaech et al., 2024) has demonstrated remarkable performance in solving reasoning tasks, sparking widespread interest across domains in test-time scaling research. By leveraging additional computation during inference time to conduct ”slow thinking” (Kirillov et al., 2023), it further enhances the accuracy of responses to complex problems.
@@
Inspired by the extensive research on the CoT in LLMs, reasoning tasks in multi-modal large language models (MLLMs) (Wang et al., 2024a; Bai et al., 2025; Team et al., 2025; Wu et al., 2024a; Maaz et al., 2023) have experienced rapid advancements. Typical methods include Best-of-N, Beam Search, and Monte Carlo Tree Search (Xu et al., 2024; Tian et al., 2023; Yao et al., 2024; Wang et al., 2022; Chen et al., 2024c). These approaches rely on intricate search mechanisms to generate large volumes of reasoning data and utilize supervised fine-tuning to enable models to acquire autonomous reasoning capabilities. With the advancement of reinforcement learning (RL) theory and technology, DeepSeek R1 (Guo et al., 2025) demonstrates how large language models can autonomously develop complex reasoning abilities through simple rule-based incentive mechanisms and lightweight reinforcement learning (GRPO (Shao et al., 2024)) algorithms. This method enables LLMs to naturally exhibit ”Aha Moment” without explicit supervision, characterized by self-reflection and spontaneous increases in response length during training. Recent studies (Huang et al., 2025; Zhou et al., 2025; Liu et al., 2025d; Pan et al., 2025) have extended this approach to MLLMs and applied it to areas such as object recognition (Liu et al., 2025d), semantic segmentation (Liu et al., 2025c), and video analysis (Sun et al., 2025). These methods significantly enhance MLLMs’ capabilities with limited training data, achieving performance comparable to supervised fine-tuning (SFT) in in-domain tests while outperforming SFT models in out-of-distribution (OOD) evaluations. However, this rapid progress, as illustrated in Figure 1, presents significant challenges to researchers. Although RL-based methods are effective, most of them adhere to text-based thinking practices, neglecting the critical role of other modalities in multimodal scenarios. Moreover, current RL-based reasoning methods primarily rely on rule-based reward functions with verifiable answers, overlooking broader general-scenario problems where verifiable answers are unavailable.
@@
Although many existing surveys focus on MLLMs reasoning (Wang et al., 2025c; Lin et al., 2025b), none of them specifically address the issue of RL-based reasoning for MLLMs. To fill this gap, we provide a comprehensive and systematic review of reasoning methods for RL-based MLLMs, offering a structured analysis of technological advancements, methodologies, practical applications, and future directions. Our goal is to provide researchers with a thorough and systematic guide to identifying appropriate methods in the rapidly evolving field of MLLM reasoning, thereby promoting further innovation and progress in this dynamic area. We first introduce the background of MLLMs, Chain-of-Thought reasoning, and reinforcement learning in Section 2. Then, we review key RL algorithm designs and optimizations for both LLMs and MLLMs in Section 3. Next, we survey RL-based reasoning methods for MLLMs, covering algorithmic strategies, reward mechanisms, and benchmark evaluations in Section 4, Section 5 and Section 6. Finally, Section 7 discuss current limitations and future research directions in this rapidly evolving field in.
@@
This paper addresses RL-based reasoning for Multimodal Large Language Models (MLLMs) from four key perspectives:
@@
  • Explore the Key Designs and Optimizations of RL in LLMs and MLLMs, with a focus on analyzing the core ideas and improvement directions of value-free methods and value-based methods. We analyze the innovative solutions these methods employ to address challenges in training efficiency, stability, and performance, and discuss their respective strengths, weaknesses, and future optimization opportunities.
  • Analysis of existing RL-based reasoning methods in terms of algorithmic frameworks, reward function designs, and modality integration strategies. We systematically categorize representative approaches based on their use of reinforcement learning algorithms, the design of accuracy- and structure-oriented reward mechanisms, and the ways in which they incorporate multimodal inputs including vision, audio, and temporal information.
  • Survey of benchmark datasets and evaluation protocols used to assess MLLMs’ reasoning capabilities. We examine the construction pipelines of aligned datasets, covering data sources, model response collection, and preference annotation methods. The surveyed benchmarks span mathematical, scientific, spatial, and interaction-based reasoning tasks, and are organized by domain specificity and generalization capacity.
@@
  • Identification of current limitations and future research directions in RL-based reasoning for MLLMs. We discuss unresolved challenges such as sparse and static reward feedback, inefficient reasoning trajectories, and weak cross-modal coordination. Promising directions include hierarchical reward modeling, vision-guided CoT generation, and lightweight RL frameworks suitable for real-world multimodal agents.
@@ 2.1. Multi-Modal Large Language Models and MM-COT
Despite the impressive zero-shot/few-shot reasoning performance of LLMs in most Natural Language Processing (NLP) tasks, they are limited to understanding only discrete textual information. In contrast, models for other modalities are capable of integrating information from various modalities such as images, audio, and video (Kirillov et al., 2023; Shen et al., 2024; Oquab et al., 2023), but often fall short in reasoning capabilities. Given this complementarity, the integration of LLMs with other specialized models has led to the emergence of multimodal large language models. Representative examples include open-source models such as Qwen-VL (Bai et al., 2023; Wang et al., 2024a; Bai et al., 2025), Intern-VL (Chen et al., 2024f, e; Zhu et al., 2025), LLaVA (Liu et al., 2023, 2024b), and Valley (Luo et al., 2023; Wu et al., 2025), as well as closed-models like GPT-4o (Hurst et al., 2024) and Gemini (Team et al., 2023, 2024). These MLLMs leverage LLMs as the core cognitive engine, demonstrating exceptional capabilities such as strong language generation, zero-shot transfer, and ICL. At the same time, foundational models for other modalities empower a variety of multimodal tasks by providing high-quality representations for non-textual data. In summary, mainstream MLLMs typically follow a consistent architecture: they process multimodal embeddings or tokens through a decoder structure, generating context-aware outputs autoregressively. Some studies have attempted to enable MLLMs to handle arbitrary modality conversions (Bai et al., 2025; Xu et al., 2025a; Hurst et al., 2024), further advancing the development of ”anything-to-anything” paradigm models.
Figure 2. Illustration of Different Modalities of Chain-of-Thought. The upper subplot shows a text-only CoT, while the lower subplot presents a multimodal CoT.
Chain-of-Thought is a technique that encourages models to generate intermediate reasoning steps $S=(s_{1},\dots,s_{n})$ before arriving at the final answer $A$. In the context of multimodal reasoning tasks, given a prompt $P$ and query $Q$, each reasoning step $s_{i}$ is sampled from the MLLM $\mathcal{M}$:
(1) $p(S|P,Q)=\prod_{i=1}^{|S|}\prod_{j=1}^{|s_{i}|}\mathcal{M}(s_{i}|P,Q,S_{<i},s_{<j})$
@@
Unlike traditional CoT, Multimodal CoT (MM-CoT) attempts to incorporate multimodal information into $P$, $S$, and $A$ to varying degrees. As illustrated in Figure 2, we categorize MM-CoT into two distinct scenarios: one that relies solely on linguistic information and another that integrates multimodal signals beyond textual content. The former aims to solve tasks involving multi-modal inputs using purely text-based CoT, while the latter emphasizes the integration of given, retrieved, or generated multi-modal information within the MM-CoT itself. Since most popular MLLMs still struggle to generate images or other modalities, recent advancements in RL-based reasoning have primarily concentrated on the generation of text-only CoT.
@@ 2.2.1. Policy Optimization Methods
Reinforcement Learning is a fundamental approach in machine learning where an agent learns to interact with an environment by taking actions, receiving rewards, and updating its policy to maximize long-term returns. Reinforcement Learning with Human Feedback (RLHF) (Bai et al., 2022), leveraging algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), guides model behavior to enhance the alignment, coherence, and utility of generated responses.
@@
In recent years, RL has increasingly been applied to enhance the reasoning capabilities of LLMs (Guo et al., 2025; Chen et al., 2024a; Team et al., 2025; Zhang et al., 2024e). By treating reasoning token generation as a Markov Decision Process, models are trained to maximize the expected return of reasoning paths, guiding optimization toward more structured and coherent reasoning trajectories. In PPO, both the policy and critic models must be trained simultaneously, which imposes significant computational demands when model parameters or token counts are large. To reduce the resource consumption of PPO while stabilizing the training process, some studies have begun exploring ways to simplify the PPO training pipeline.
@@ 2.2.2. Reward Mechanisms
In contrast to ORM, PRM place emphasis on evaluating the model’s intermediate behaviors during the reasoning process, thereby encouraging the incremental formation of coherent and logically sound CoT. Rather than focusing solely on task completion, PRM rewards intermediate behaviors that align with desirable reasoning patterns, even in cases where the final objective is not fully achieved. In model inference tasks, this approach is commonly employed to promote the generation of logically consistent intermediate reasoning steps, plausible deductive structures, or the use of credible sources from external knowledge bases. Recent research has proposed the incorporation of process-level reward signals—such as logical consistency, informational completeness, or citation reliability for each reasoning step—into the fine-tuning process, thereby improving model interpretability and enhancing user trust (Andrychowicz et al., 2020; Lightman et al., 2023; Baker et al., 2025). Compared to outcome-based rewards, process-level mechanisms offer finer-grained supervision, facilitating more effective debugging and alignment of model behavior. Nevertheless, the design of process rewards depends heavily on the accurate evaluation of intermediate reasoning steps, which poses significant challenges for automation. Moreover, the absence of standardized and robust evaluation criteria limits the scalability and generalizability of PRMs across diverse tasks and model architectures.
@@ 2.2.3. Training Efficiency
Data-efficient learning aims to maximize performance with minimal data. Two commonly adopted approaches in this category include: (1) Prioritized Sampling (Team et al., 2025): Rather than uniformly sampling from experience replay buffers, prioritized sampling selects transitions based on their learning potential. This strategy focuses updates on informative samples, accelerating learning and improving policy quality. (2) High-Quality Samples (Wen et al., 2025): Leveraging samples with higher utility—such as trajectories with higher rewards or successful episodes—can significantly improve sample efficiency. By concentrating training on more informative experiences, the agent learns more effectively while reducing unnecessary computational overhead.
@@ 2.3.1. Differences and Similarities of Reasoning in LLMs and MLLMs
LLMs and MLLMs both employ CoT-based reasoning to address complex tasks, with the core principle being the decomposition of intricate problems into intermediate steps that guide the model toward a solution. Whether prompted through textual examples (for LLMs) or multimodal inputs (for MLLMs), the CoT process remains fundamentally similar: the model systematically builds its reasoning based on prior steps, providing transparency into its internal processes. This transparency is particularly valuable for tasks like mathematical reasoning and code generation, where intermediate steps are crucial for interpreting outputs. However, significant differences arise due to the types of inputs processed. We summarize the main differences in CoT-based reasoning between LLMs and MLLMs in Table 1. In LLMs, CoT reasoning is confined to linguistic operations, involving logical deduction, arithmetic computation, or language clarification within a linear, text-centric framework. Conversely, MLLMs engage in more dynamic and nonlinear reasoning by alternating between textual and other modalities, such as visual perception. For instance, an MLLM might first reason about visual input (e.g., identifying ”a blue car” in an image) and then integrate this with textual reasoning (e.g., referencing typical car colors). This integration of modalities makes MLLMs’ CoT reasoning inherently multidimensional but introduces challenges like cross-modal alignment, accurate incorporation of visual features, and avoiding modality bias. Grounding further differentiates the two, as MLLMs anchor reasoning in real-world knowledge, such as visual objects or spatial relationships, while LLMs rely solely on textual and abstract concepts. In summary, while both models leverage CoT reasoning, LLMs excel in text-centered reasoning, whereas MLLMs expand this capability by integrating multimodal data, necessitating tailored approaches that account for their unique characteristics and enable nonlinear interactions across modalities in MM-CoT reasoning.

### v2

@@ Abstract.
The application of reinforcement learning (RL) to enhance the reasoning capabilities of Multimodal Large Language Models (MLLMs) constitutes a rapidly advancing research area. While MLLMs extend Large Language Models (LLMs) to handle diverse modalities such as vision, audio, and video, enabling robust reasoning across multimodal inputs remains challenging. This paper provides a systematic review of recent advances in RL-based reasoning for MLLMs, covering key algorithmic designs, reward mechanism innovations, and practical applications. We highlight two main RL paradigms, value-model-free and value-model-based methods, and analyze how RL enhances reasoning abilities by optimizing reasoning trajectories and aligning multimodal information. Additionally, we provide an extensive overview of benchmark datasets, evaluation protocols, and current limitations, and propose future research directions to address challenges such as sparse rewards, inefficient cross-modal reasoning, and real-world deployment constraints. Our goal is to provide a comprehensive and structured guide to RL-based multimodal reasoning.
@@ 1. Introduction
The emergence of large language models (LLMs) (Yang et al., 2024; GLM et al., 2024; Grattafiori et al., 2024) has ushered in an unprecedented new era for the field of artificial intelligence, showcasing exceptional capabilities such as instruction following and few-shot learning (Brown et al., 2020). However, achieving human-level intelligence requires not only surpassing basic perceptual abilities but also developing sophisticated cognitive reasoning skills that enable iterative reasoning through contextual understanding and self-correction. Inspired by this, in-context learning (ICL) techniques (Wei et al., 2022; Wu et al., 2022) have empowered LLMs with step-by-step reasoning abilities, commonly referred to as the Chain-of-Thought (CoT) reasoning mechanism (Wang and Zhou, 2024; Besta et al., 2024). OpenAI’s o1 model (Jaech et al., 2024) has demonstrated remarkable performance in solving reasoning tasks, sparking widespread interest across domains in test-time scaling research. By leveraging additional computation during inference time to conduct ”slow thinking” (Kirillov et al., 2023), it further enhances the accuracy of responses to complex problems.
@@
Inspired by the extensive research on the CoT in LLMs, reasoning tasks in multi-modal large language models (MLLMs) (Wang et al., 2024a; Bai et al., 2025; Team et al., 2025a) have experienced rapid advancements. Typical methods include Best-of-N, Beam Search, and Monte Carlo Tree Search (Xu et al., 2024; Yao et al., 2024; Chen et al., 2024b). These approaches rely on intricate search mechanisms to generate large volumes of reasoning data and utilize supervised fine-tuning (SFT) to enable models to acquire autonomous reasoning capabilities. With the advancement of reinforcement learning (RL) theory and technology, DeepSeek R1 (Guo et al., 2025) demonstrates how large language models can autonomously develop complex reasoning abilities through simple rule-based incentive mechanisms and lightweight reinforcement learning (GRPO (Shao et al., 2024)) algorithms. This method enables LLMs to naturally exhibit ”Aha Moment” without explicit supervision, characterized by self-reflection and spontaneous increases in response length during training. Recent studies (Huang et al., 2025; Zhou et al., 2025a; Liu et al., 2025f; Pan et al., 2025) have extended this approach to MLLMs and applied it to areas such as object recognition (Liu et al., 2025f), semantic segmentation (Liu et al., 2025e), and video analysis (Sun et al., 2025). These methods significantly enhance MLLMs’ capabilities with limited training data, achieving performance comparable to SFT in in-domain tests while outperforming SFT models in out-of-distribution (OOD) evaluations. However, this rapid progress, as illustrated in Figure 1, presents significant challenges to researchers. Although RL-based methods are effective, most of them adhere to text-based thinking practices, neglecting the critical role of other modalities in multimodal scenarios. Moreover, current RL-based reasoning methods primarily rely on rule-based reward functions with verifiable answers, overlooking broader general-scenario problems where verifiable answers are unavailable.
@@
Although many existing surveys focus on MLLMs reasoning (Wang et al., 2025d; Lin et al., 2025b), none of them specifically address the issue of RL-based reasoning for MLLMs. To fill this gap, we provide a comprehensive and systematic review of reasoning methods for RL-based MLLMs, offering a structured analysis of technological advancements, methodologies, practical applications, and future directions. Our goal is to provide researchers with a thorough and systematic guide to identifying appropriate methods in the rapidly evolving field of MLLM reasoning, thereby promoting further innovation and progress in this dynamic area. We first introduce the background of MLLMs, CoT reasoning, and reinforcement learning in Section 2. Then, we review key RL algorithm designs and optimizations for both LLMs and MLLMs in Section 3. Next, we survey RL-based reasoning methods for MLLMs, covering algorithmic strategies, reward mechanisms, and benchmark evaluations in Section 4, Section 5 and Section 6. Finally, Section 7 discuss current limitations and future research directions in this rapidly evolving field in.
@@
This paper addresses RL-based reasoning for MLLMs from four key perspectives:
@@
  • Core Designs of RL in LLMs/MLLMs : Analyze value-model-free and value-model-based methods, focusing on their innovations in training efficiency, stability, and performance, while discussing strengths, weaknesses, and optimization opportunities.
  • Analysis of RL-Based Reasoning Methods : Examine the algorithmic frameworks, the design of reward functions including both accuracy-oriented and structure-oriented approaches, and the strategies for integrating multiple modalities. Furthermore, emphasize how each of these components contributes to addressing the core challenges of multimodal reasoning.
  • Benchmark Datasets and Evaluation Protocols : Survey datasets and evaluation methods for reasoning tasks, covering construction pipelines such as data sources and annotations, as well as benchmarks spanning math, science, spatial, and interaction-based domains.
@@
  • Limitations and Future Directions : Highlight unresolved challenges such as sparse rewards, inefficient trajectories, and cross-modal coordination. Meanwhile, propose future directions including hierarchical reward modeling, vision-guided CoT, and lightweight RL frameworks for real-world use.
@@ 2.1. Multi-Modal Large Language Models and MM-COT
Despite the impressive zero-shot/few-shot reasoning performance of LLMs in most Natural Language Processing (NLP) tasks, they are limited to understanding only discrete textual information. In contrast, models for other modalities are capable of integrating information from various modalities such as images, audio, and video (Bai et al., 2025; Li et al., 2025b; Sun et al., 2025), but often fall short in reasoning capabilities. Given this complementarity, the integration of LLMs with other specialized models has led to the emergence of multimodal large language models. Representative examples include open-source models such as Qwen-VL (Wang et al., 2024a; Bai et al., 2025), Intern-VL (Chen et al., 2024d; Zhu et al., 2025), LLaVA (Liu et al., 2023, 2024b), and Valley (Luo et al., 2023; Wu et al., 2025), as well as closed-models like GPT-4o (Hurst et al., 2024) and Gemini (Team et al., 2023, 2024). These MLLMs leverage LLMs as the core cognitive engine, demonstrating exceptional capabilities such as strong language generation, zero-shot transfer, and ICL. At the same time, foundational models for other modalities empower a variety of multimodal tasks by providing high-quality representations for non-textual data. In summary, mainstream MLLMs typically follow a consistent architecture: they process multimodal embeddings or tokens through a decoder structure, generating context-aware outputs autoregressively. Some studies have attempted to enable MLLMs to handle arbitrary modality conversions (Bai et al., 2025; Xu et al., 2025a), further advancing the development of ”anything-to-anything” paradigm models.
Figure 2. Illustration of Different Modalities of CoT. The upper subplot shows a text-only CoT, while the lower subplot presents a multimodal CoT.
CoT is a technique that encourages models to generate intermediate reasoning steps $S=(s_{1},\dots,s_{n})$ before arriving at the final answer $A$. In the context of multimodal reasoning tasks, given a prompt $P$ and query $Q$, each reasoning step $s_{i}$ is sampled from the MLLM $\mathcal{M}$:
(1) $p(S|P,Q)=\prod_{i=1}^{|S|}\prod_{j=1}^{|s_{i}|}\mathcal{M}(s_{i}|P,Q,S_{<i},s_{i<j})$
@@
Unlike traditional CoT, Multimodal Chain-of-Thought (MM-CoT) attempts to incorporate multimodal information into $P$, $S$, and $A$ to varying degrees. As illustrated in Figure 2, we categorize MM-CoT into two distinct scenarios: one that relies solely on linguistic information and another that integrates multimodal signals beyond textual content. The former aims to solve tasks involving multi-modal inputs using purely text-based CoT, while the latter emphasizes the integration of given, retrieved, or generated multi-modal information within the MM-CoT itself. Since most popular MLLMs still struggle to generate images or other modalities, recent advancements in RL-based reasoning have primarily concentrated on the generation of text-only CoT.
@@ 2.2.1. Policy Optimization Methods
Reinforcement Learning is a fundamental approach in machine learning where an agent learns to interact with an environment by taking actions, receiving rewards, and updating its policy to maximize long-term returns. Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022), leveraging algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), guides model behavior to enhance the alignment, coherence, and utility of generated responses.
@@
In recent years, RL has increasingly been applied to enhance the reasoning capabilities of LLMs (Guo et al., 2025; Team et al., 2025a). By treating reasoning token generation as a Markov Decision Process, models are trained to maximize the expected return of reasoning paths, guiding optimization toward more structured and coherent reasoning trajectories. In PPO, both the policy and critic models must be trained simultaneously, which imposes significant computational demands when model parameters or token counts are large. To reduce the resource consumption of PPO while stabilizing the training process, some studies have begun exploring ways to simplify the PPO training pipeline.
@@ 2.2.2. Reward Mechanisms
In contrast to ORM, PRM place emphasis on evaluating the model’s intermediate behaviors during the reasoning process, thereby encouraging the incremental formation of coherent and logically sound CoT. Rather than focusing solely on task completion, PRM rewards intermediate behaviors that align with desirable reasoning patterns, even in cases where the final objective is not fully achieved. In model inference tasks, this approach is commonly employed to promote the generation of logically consistent intermediate reasoning steps, plausible deductive structures, or the use of credible sources from external knowledge bases. Recent research has proposed the incorporation of process-level reward signals—such as logical consistency, informational completeness, or citation reliability for each reasoning step—into the fine-tuning process, thereby improving model interpretability and enhancing user trust (Zhang et al., 2025a; Fan et al., 2025). Compared to outcome-based rewards, process-level mechanisms offer finer-grained supervision, facilitating more effective debugging and alignment of model behavior. Nevertheless, the design of process rewards depends heavily on the accurate evaluation of intermediate reasoning steps, which poses significant challenges for automation. Moreover, the absence of standardized and robust evaluation criteria limits the scalability and generalizability of PRMs across diverse tasks and model architectures.
@@ 2.2.3. Training Efficiency
Data-efficient learning aims to maximize performance with minimal data. Two commonly adopted approaches in this category include: (1) Prioritized Sampling (Team et al., 2025a): Rather than uniformly sampling from experience replay buffers, prioritized sampling selects transitions based on their learning potential. This strategy focuses updates on informative samples, accelerating learning and improving policy quality. (2) High-Quality Samples (Kang et al., 2025): Leveraging samples with higher utility—such as trajectories with higher rewards or successful episodes—can significantly improve sample efficiency. By concentrating training on more informative experiences, the agent learns more effectively while reducing unnecessary computational overhead.
@@ 2.3.1. Differences and Similarities of Reasoning in LLMs and MLLMs
LLMs and MLLMs both employ CoT-based reasoning to address complex tasks, with the core principle being the decomposition of intricate problems into intermediate steps that guide the model toward a solution. Whether prompted through textual examples for LLMs or multimodal inputs for MLLMs, the CoT process remains fundamentally similar: the model systematically builds its reasoning based on prior steps, providing transparency into its internal processes. This transparency is particularly valuable for tasks like mathematical reasoning and code generation, where intermediate steps are crucial for interpreting outputs. However, significant differences arise due to the types of inputs processed. We summarize the main differences in CoT-based reasoning between LLMs and MLLMs in Table 1. In LLMs, CoT reasoning is confined to linguistic operations, involving logical deduction, arithmetic computation, or language clarification within a linear, text-centric framework. Conversely, MLLMs engage in more dynamic and nonlinear reasoning by alternating between textual and other modalities, such as visual perception. For instance, an MLLM might first reason about visual input such as identifying ”a blue car” in an image, and then integrate this with textual reasoning, such as referencing typical car colors. This integration of modalities makes MLLMs’ CoT reasoning inherently multidimensional but introduces challenges like cross-modal alignment, accurate incorporation of visual features, and avoiding modality bias. Grounding further differentiates the two, as MLLMs anchor reasoning in real-world knowledge, such as visual objects or spatial relationships, while LLMs rely solely on textual and abstract concepts. In summary, while both models leverage CoT reasoning, LLMs excel in text-centered reasoning, whereas MLLMs expand this capability by integrating multimodal data, necessitating tailored approaches that account for their unique characteristics and enable nonlinear interactions across modalities in MM-CoT reasoning.